OpenStudy (anonymous):

At t= 0, a particle starts from rest at x= 0, y = 0, and moves in the xy plane with an acceleration a⃗ = (4.0i^ + 3.0j^)m/s2. Determine the position of the particle

OpenStudy (anonymous):

you can use double integrals to solve this problem \[r - ro =\int\limits\limits_{}^{}\int\limits\limits_{}^{} (4 i + 3 j )dt\]

OpenStudy (anonymous):

make sure it's dt^2 ^_^ \[ r-r_o = \iint(4 \ \hat{ \textbf i} + 3 \ \hat{ \textbf j}) dt^2 \]

OpenStudy (anonymous):

Ah, ya. I missed it. Thanks :)